OpenStudy (anonymous):

Solve 152x = 36. Round to the nearest ten-thousandth. A. 0.6616 B. 2.6466 C. 1.7509 D. 1.9091

pooja195 (pooja195):

152x=36 divide both sides by 152 x=36/152 36/152=?

OpenStudy (anonymous):

i got .236 ? @pooja195

pooja195 (pooja195):

yep

OpenStudy (anonymous):

but that not an answer

pooja195 (pooja195):

hold on a sec u did put 152 first right

OpenStudy (anonymous):

no i just did it and got 4.22 ? still not an option

pooja195 (pooja195):

wat r yr options??

OpenStudy (anonymous):

A. 0.6616 B. 2.6466 C. 1.7509 D. 1.9091

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

hold on

pooja195 (pooja195):

I think something is wrong with what you posted. To solve, you need to get x alone so you have to divide both sides by 152 and the answer that I get is .2368

OpenStudy (anonymous):

oh hold on 15^2x=36 sorry

OpenStudy (anonymous):

the 2x is raised to 15 sorry

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

use logarithms 15^(2x) = 36 2xlog15= log36 x= log36/(2* log15) x= aprox 0.6616

OpenStudy (anonymous):

okay thank you sorry

pooja195 (pooja195):

np